OpenStudy (anonymous):

I NEED HELP! Use the quadratic formula to solve the equation: 6x 2 − 5x + 1 = 0 A. x = 1 or x = – 1 ⁄ 2 B. x = 2 or x = −1 C. x = – 1 ⁄ 3 or x = 3 D. x = 1 ⁄ 2 or x = 1 ⁄ 3

OpenStudy (anonymous):

the quadratic formula is \[\large x = \frac{ -b \pm \sqrt{b^2 - 4ac}}{2a}\] if you substitute you will get \[\large x = \frac{-(-5) \pm \sqrt{(-5)^2 - 4(6)(1)}}{2(6)}\]

OpenStudy (anonymous):

therefore \[\large x = \frac{5 \pm \sqrt{25 - 24}}{12}\] \[\large x = \frac{5 \pm 1}{12}\] \[\large x = \frac{4}{12}; \frac{6}{12}\]

OpenStudy (anonymous):

that cant be... uhhh

OpenStudy (anonymous):

\[\large x = \frac{1}{3}; \frac{1}{2}\]

OpenStudy (anonymous):

Thank you